Activity at the Soufriere Hills volcano has increased over the last week. A major new dome with a permanent spine first appeared 15 December; the spine has grown continuously since then at a vertical rate of up to 20 m/day. Seismic activity has continued at a low rate, with occasional large long-period earthquakes occurring more frequently than before. Deformation measurements showed a sudden change on 19 December. EDM measurements have been restricted by persistent low cloud, but in general have shown only small changes.
Seismic activity has continued at a similar rate to last week. Locateable earthquakes have occurred at a rate of 2 to 20 per day. Almost all the volcano-tectonic earthquakes have been located away from the volcano, in particular at St George's Hill and in the Long Ground area. Depths have been in the range of 0 to 6 km. The small, frequent long-period earthquakes recorded by the Gages seismometer have continued intermittently, but have not reached the size or rate observed before 9 December. The number of large long-period earthquakes has increased to a rate of 5 to 15 per day. These events have a dominant frequency of 2.2 Hz, and emergent onsets, and are recorded by all stations. Some of these earthquakes have been accompanied by small ash clouds that are visible from the observatory. This suggests that these earthquakes are caused by small explosions within the crater. Heavy rain from 16 to 19 December caused many floods and dilute mudflows around the volcano, which were recorded by the seismic network.
EDM measurements were restricted this week by the low cloud. The lines to the Castle Peak target were measured three times, and showed continued slow outward movement of Castle Peak, at a rate of about 1 cm/day. The Gages Wall target was measured for the first time on 17 December, and remeasured on 19 and 20 December. An increase in line length of 4 cm was detected on 19 December. Measurement of the other EDM lines showed no significant changes.
GPS sites at St George's Hill, Harris' Lookout and Reid's Hill were occupied each day. The Radio Antilles site was relocated to Reid's Hill on 14 December. Measurements up to 18 December showed essentially no change in line lengths. Data recorded on 18-19 December gave a 4 cm lengthening of both the lines to Harris' Lookout. This change was coincident with the movement of the Gages EDM target, and was the largest single change recorded by GPS during this crisis. On the following day, the Reid's Hill to Harris' Lookout length decreased by 1.5 cm (the St George's station did not record data due to a power failure). Occupation of the GPS network will continue until 28 December, although the data will no longer be processed immediately. Permanent installation of a two-station GPS network is planned for mid-January.
The two electronic tiltmeters continued to show no significant changes.
The EDM and GPS deformation measurements suggest continued slight inflation of the volcano. The larger changes of the GPS lines to Harris's Lookout on 18-19 December suggest there may have been a temporary increase in inflation. However the sign of the Gages Wall EDM line change is not conisistent with inflation, and the higher rate of movement has not continued.
Good visual observations were possible only on 15, 17 and 20 December. Observations from the helicopter on 15 December confirmed the presence of a new area of dome growth in the west of the crater, which had spalled material into the moat. A prominent new spine was first observed on 17 December, at the eastern side of the crater, just south of the spine that grew between 2 and 10 December. This spine had an estimated height of 40 m on 17 December. Accurate measurements on 17 and 20 December gave a daily growth rate of 7 m/day for the spine, and a slightly lower rate for the top of the dome.
Observers in the helicopter on 17 December saw increased glowing from all areas of the new lava dome. The dome in the July 18th vent had a narrow crack that was emitting small glowing rocks continuously, and a large hot rock fell from the top of the crack into the moat. Glowing rock is now clearly visible from the Long Ground/Whites area and the airport. Small hot rock avalanches have been reported, but no material has yet left the crater area. Many small ash releases have occurred, with ash clouds up to 2000 m above sea level.
The active dome growth since about 13 December has been nearly aseismic unlike previous episodes of dome growth. This suggests that magma now has a more free path to the surface.